when methane gas is mixed with chlorine gas, hydrochloric acid is produced. the balanced equation is: CH4(g) + 2(Cl2)(g) ->4HCl(g) +CCl4(g) What mass of hydrochloric acid is produced if 1.25 g of CH4 reacts in excess of chlorine gas?


1
Expert's answer
2021-09-14T01:24:02-0400

Mole ratio of CH4 :HCl = 1:4

Molar mass of CH4= 16.04 g/mol

1.25/16.04 = 0.078 mol

Molar mass of HCl = 36.458 g/mol

= 0.078 × 4 × 36.458 = 11.365g